OpenStudy (jokersmart):

Describe the best way to solve 27 x 4 using mental math.

OpenStudy (babylove18):

4 times 20 then 4 times 7 and add them together

imqwerty (imqwerty):

^

OpenStudy (babylove18):

4*20= 80 + 4*7 = 28 so 80 + 28 = 108